Express the number appearing in the following statements in standard form.
 Size of bacteria is 0.0000005 m

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Size of bacteria = 0.0000005
= 5/10000000 = 5/10^7= 5×10^-7m

Thanks